Find a1 if d = 4, a20 = 82, and S20 = 880.
A.) 2
B.) 4
C.) 6
D.) 60
is this the right rule?
an = a1+(n-1)d
Sn = (n/2)(a1+an)
If so can I see a demonstration on how you solve it?

Found 2 solutions by stanbon, ikleyn:
Answer by stanbon(75887)

Find a1 if d = 4, a20 = 82, and S20 = 880.
A.) 2
B.) 4
C.) 6
D.) 60
----------------------
an = a1+(n-1)d
Sn = (n/2)(a1+an)
----------
a(20) = a(1) + 19*4 = a(1) + 76 = 82
a(1) = 82-76 = 6
------------------
S(20) = (20/2)(a(1) + a(20)) = 10(a(1) + 82)) = 880
a(1) + 82 = 88
a(1) = 6
----------------------
Cheers,
Stan H.
==================

Answer by ikleyn(52781)

.
On arithmetic progressions see the lesson
    - Arithmetic progressions
in this site.